OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−2, 0) and (2, 8)?

OpenStudy (anonymous):

Find the gradient. Then put the gradient in the equation(choosing 2,8): y-8=m(x-2) That is the equation.

OpenStudy (jdoe0001):

hmmm what' s the slope of those 2 points?

OpenStudy (anonymous):

\[m=\Delta X/\Delta Y\]\[m=(8-0)/(2-(-2))\]\[m=8/4\]\[m=2\]

OpenStudy (anonymous):

choices y = 2(x + 2) y = 2(x − 2) y = −2(x − 2) y = −2(x + 2)

OpenStudy (jdoe0001):

hmmm that doesn't look like the Point-Slope form

OpenStudy (anonymous):

It should be y=2(x+2), after putting m=2 into the equation above.